3x+2y=20
-2y+x=-6y

Answer by stanbon(75887)

3x+2y=20
-2y+x=-6y
----------------
Rearrange::
3x + 2y = 60
x = -4y
------
Substitute for "x" and solve for "y"::
3(-4y)+2y = 60
-12y+2y = 60
-10y = 60
y = -6
-----
x = -4y = -4(-6) = 24
------------
Cheers,
Stan H.
---------